OpenStudy (vshiroky):

Fina an equation of the line that has the given slope and passes through the given point. m=-3 (1/2,2)

OpenStudy (polaris_s0i):

\[ y = -3x + b \] \[ 2 = -3x + \frac{1}{2} \] now solve.

OpenStudy (vshiroky):

I got y=-3x+5/2

OpenStudy (polaris_s0i):

er sorry \[ 2 = -3(\frac{1}{2}) + b \]

OpenStudy (vshiroky):

is mine correct?

OpenStudy (polaris_s0i):

you tell me... plugin (1/2) into x

OpenStudy (vshiroky):

Isn't it y-y1=-3(x-x1)

OpenStudy (vshiroky):

I did the problem I just don't know if I am correct

OpenStudy (polaris_s0i):

no, you're confused... we aren't finding the distance of anything, no y1 y2 stuff here. just set x to 1/2 and solve for y, if you get 2 then you are correct.

OpenStudy (vshiroky):

Yeah I dont get it

OpenStudy (polaris_s0i):

lol, maybe I'll explain then.

OpenStudy (vshiroky):

It's supposed to be in slope intercept form

OpenStudy (vshiroky):

Is it y=-3x+9/2

OpenStudy (polaris_s0i):

line has slope -3. we know a line has an equation of the form y = mx +b where m is the slope. so y = -3x + b is what we know. We also know the point it's going through, so that is a specific x and y coordinate. we can set the equation to: \[ 2 = -3(\frac{1}{2}) + b\] and solve for b, because we absolutely want the line to go through this point. \[ 2 = -\frac{3}{2} + b \] \[ \frac{4}{2} + \frac{3}{2} = b \] \[ \frac{7}{2} = b \] so now we have the equation \[ y = -3x + \frac{7}{2} \] if we plug our original x coordinate in, we had better get 2, since that is what we designed it to do: \[ y = -3\frac{1}{2} + \frac{7}{2} \] \[ y = -\frac{3}{2} + \frac{7}{2} \] \[ y = \frac{4}{2} \] \[ y = 2 \]

OpenStudy (vshiroky):

This isn't slope intercept form though

OpenStudy (polaris_s0i):

"slope-intercept" form: y = mx + b

OpenStudy (vshiroky):

Right so wouldn't it be y=-3x+9/2

OpenStudy (vshiroky):

You just have y=2

OpenStudy (polaris_s0i):

where is the 9/2 coming from? I have: \[ y=−3x+\frac{7}{2} \]

OpenStudy (polaris_s0i):

y =2 was me checking to make sure its write.

OpenStudy (vshiroky):

When I added the 2 from the other side

OpenStudy (polaris_s0i):

can you show me you're calculations?

OpenStudy (vshiroky):

I am running out of time

OpenStudy (vshiroky):

I did y-2=-3(x-1/2) y-2=-3x+7/2 y=-3x+9/2

OpenStudy (polaris_s0i):

no, that is incorrect. you're not doing the correct thing. you need to READ THE WHOLE ANSWER I gave above, because you are not understanding the WHY of what you are doing, you're following method without reason.

OpenStudy (vshiroky):

Ok